OpenStudy (anonymous):

Which is an equation in slope-intercept form of the line that contains the points S(7, –9) and T(8, –7)?

OpenStudy (anonymous):

The slope is \[\frac{-7-(-9)}{8-7}=2\]So we have \[y=2x+b\]where b is the y-int (which we dont have yet). The y-int is the point where x=0, (0,b). Since we know the slope (rise over run=2) we can get the value of b:\[2=\frac{-9-b}{7-0}\]Solving...\[b=-23\]So our line is:\[y=2x-23\]